A Complete Bibliography of Games: Research and Practice

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 August 2024 Version 1.01

Title word cross-reference

3 [12].

across [15]. Action [38]. Activities [29]. Advancing [46]. Advertising [33]. Affective [19]. AI [44, 14]. Analysis [4]. Applying [40]. Asia [27]. Ask [45]. Audio [16, 12]. Avoidance [38].

Back [10]. Based [34, 33]. Been [32]. Behavioral [43]. Belgian [42]. Benefits [12]. Best [32]. Beyond [16, 13]. Box [25]. Broaden [22].

C [22]. Call [38]. Case [33]. Casual [15]. Cataclysmic [28]. Causes [37]. centric [45]. Challenges [3, 29]. Charting [13]. Chat [21]. Cheap [40]. Cheating [39]. Children [35]. Citizens [14]. Co [31]. Co-Creating [31]. Cognition [15]. Combat [24]. Communities [39]. Compassion [32]. Competitive [40]. Compliance [25]. Concepts [36]. Conscious [3].

Content [32, 12]. Control [24]. Cost [40]. Counter [29]. Counter-extremism [29]. Creating [31]. creative [12]. Critical [15]. Cross [39]. Cross-Industry [39]. Cultures [18].

D [12]. Data [13]. Decentralized [17]. Deep [31]. Design [35, 9, 19, 3, 31, 43]. Developer [32]. Developers [34]. development [42]. Dialogue [38]. Diegesis [20]. Digital [35, 29]. Directions [36]. Diversity [4]. Do [28]. Dynamics [39].

East [27]. Editorial [1, 14]. Education [26, 42]. Embedding [35]. Embracing [8]. Emotion [19]. Environmentally [3]. Environments [41]. Epic [25]. Ethical [34, 45]. Ethics [46, 42]. Eudaimonic [20]. Evidence [34]. Evidence-Based [34]. Exposure [24]. Extremism [23, 38, 30, 29]. Extremist [29]. Extremity [4].

Faces [4, 42]. First [14]. Framework [40, 39, 19]. Friedman [40]. Fusion [18]. Future [10, 17, 11, 13]. Futures [5, 31].

Game [40, 39, 28, 3, 16, 31, 13, 41, 42]. Gamer [18]. Games [2, 46, 19, 5, 1, 33, 24, 26, 10, 34, 17, 11, 14, 21, 37, 15, 36, 7, 6, 8, 4, 43, 12]. Gaming [44, 23, 38, 12, 29, 27, 30]. Global [26]. Go [28, 17, 33]. Guidance [34]. guidelines [45]. Guiding [39].

Harassment [38]. Harm [39]. Hate [38, 41]. Health [37]. Here [28]. Host [14].

Identity [18]. Illuminated [10]. Impact [10]. implementation [12]. Inaugural [1]. Individual [24]. Industry [46, 44, 39, 23, 28, 38]. Infrastructure [8]. Input [20]. Inspired [36]. Interaction [20]. Interactivity [19]. Intervention [39].

Justice [36].

Labels [25]. Landscape [27]. Layoffs [28]. Lens [31]. Lessons [3]. Let [17]. Lifespan [15]. Lighthouse [1]. Location [33]. Location-based [33]. Locative [33]. Looking [10]. Loot [25].

Manifesto [2]. Many [42, 4]. Massive [4]. Massive-scale [4]. Matter [6]. Matters [32]. May [14]. Meaning [20]. Meaningful [23]. Media [1]. Mental [37]. Metaverse [11, 16, 21]. Microsoft [25]. microtransactions [45]. Mobile [4]. Moderating [21]. Monetisation [4]. Multiplayer [41]. Museum [36]. Museum-Inspired [36].

names [42]. Needs [44, 22]. Networks [40]. Next [12]. Nexus [30].

Nintendo [25].

Observations [27]. Off [41]. Online [23, 11, 41]. Open [38]. Opportunities [29]. Overview [36].

pitfalls [12]. Platform [33]. Platforms [7, 25]. Play [18, 15]. Playable [1]. Player [20, 4, 45]. Player-centric [45]. Players [34, 45]. Playful [35]. point [12]. Pokémon [33]. Policy [36]. Practice [9]. Practices [32]. Presence [25]. principles [45]. Probable [10]. Productive [2]. Psychological [2].

Re [40]. Re-Applying [40]. Reach [22]. Realism [12]. Reflection [20]. Reflections [27]. Research [2, 11, 36]. Resnick [40]. Responsible [44]. Rightful [14]. Rights [35]. Risks [18]. Round [46].

Safeguarding [34]. Safety [21]. scale [4]. Science [13]. Shopping [25]. Social [9, 10, 21, 7, 8]. Sociality [8]. Sony [25]. South [27]. Spaces [29]. Spending [4]. Standard [22]. Step [12]. Study [41]. Symbolic [36]. Symptoms [37]. Synthesis [15]. Systems [9].

Table [46]. Teaching [3]. technical [12]. Therapeutic [37]. Thriving [17]. Toxic [40, 18]. Toxicity [24, 41]. Transactional [4]. Transformative [31]. Transitional [36]. Transparency [23]. Treating [37].

Understanding [4]. Unpacking [39]. Unsatisfactory [25]. Usernames [40].

via [4]. Video [40, 19, 16, 37, 36, 8, 43]. view [12]. Voice [21].

Warned [32]. Warning [25]. Warnings [32]. Where [28]. World [35, 11].

You’ve [32].

References

Deterding:2023:IEL

[1] Sebastian Deterding, Kenny Mitchell, Rachel Kowert, and Brad King. Inaugural editorial: a lighthouse for games and playable media. Games: Research and Practice, 1(1):1:1–1:??, March 2023. CODEN ???? ISSN 2832-

    5516 (electronic). URL https://dl.acm.org/doi/10.1145/3585393.

Ballou:2023:MMP

[2] Nick Ballou. A manifesto for more productive psychological games research. Games: Research and Practice, 1(1):2:1–2:??, March 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3582929.

Fizek:2023:TEC

[3] Sonia Fizek, Mikhail Fiadotau, Hanna Wirman, and Maria Garda. Teaching environmentally conscious game design: Lessons and challenges. Games: Research and Practice, 1(1):3:1–3:??, March 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3583058.

Zendle:2023:MFM

[4] David Zendle, Catherine Flick, Sebastian Deterding, Joe Cutting, Elena Gordon-Petrovskaya, and Anders Drachen. The many faces of monetisation: Understanding the diversity and extremity of player spending in mobile games via massive-scale transactional analysis. Games: Research and Practice, 1(1):4:1–4:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3582927.

Deterding:2023:GF

[5] Sebastian Deterding, Kenny Mitchell, Rachel Kowert, and Brad King. Games futures I. Games: Research and Practice, 1(1):5:1–5:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3585394.

Taylor:2023:GM

[6] T. L. Taylor. Games matter. Games: Research and Practice, 1(1):6:1–6:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://

dl.acm.org/doi/10.1145/3582928.

Steinkuehler:2023:GSP

[7] Constance Steinkuehler. Games as social platforms. Games: Research and Practice, 1(1):7:1–7:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3582930.

Voll:2023:SVG

[8] Kimberly D. Voll. The sociality of video games: Embracing games as social infrastructure. Games: Research and Practice, 1(1):8:1–8:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3582931.

Cook:2023:DPS

[9] Daniel Cook. A design practice for social systems. Games: Research and Practice, 1(1):9:1–9:??, March 2023. CODEN ???? ISSN 2832-5516 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3583184.

Grace:2023:SIG

[10] Lindsay D. Grace. Social impact games, a probable future illuminated by looking back. Games: Research and Practice, 1(1):10:1–10:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3583986.

Koster:2023:OWM

[11] Raph Koster. From online world to metaverse: The future of online games and games research. Games: Research and Practice, 1(1):11:1–11:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3582932.

Rieger:2023:ANS

[12] Martin Rieger and Dominik Zingler. 3D audio: The next step in realism for gaming: Benefits and pitfalls of 3D audio in games from a creative content and technical implementation point of view. Games: Research and Practice, 1(1):12:1–12:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3583059.

Wallner:2023:BGC

[13] Günter Wallner and Anders Drachen. Beyond the game: Charting the future of game data science. Games: Research and Practice, 1(1):13:1– 13:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3582933.

Mitchell:2023:EGM

[14] Kenny Mitchell. Editorial: Games may host the first rightful AI citizens. Games: Research and Practice, 1(2):18:1–18:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3606834.

Squire:2023:CGC

[15] Kurt Squire, Garrison Wells, Maria J. Anderson-Coto, and Constance Steinkuehler. Casual games, cognition, and play across the lifespan: a critical synthesis. Games: Research and Practice, 1(2):14:1–14:??, June 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3594534.

McGregor:2023:VGA

[16] Iain McGregor. Video game audio in the metaverse and beyond. Games: Research and Practice, 1(2):15:1–15:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3593242.

Kajioka:2023:GLG

[17] E. Drake Kajioka. Games that let go: Thriving in a decentralized future. Games: Research and Practice, 1(2):16:1–16:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3600235.

Kowert:2023:YWY

[18] Rachel Kowert, Alexi Martel, and William B. Swann. You are what you play: The risks of identity fusion in toxic gamer cultures. Games: Research and Practice, 1(2):17:1–17:??, June 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3604402.

Croissant:2023:EDV

[19] Maximilian Croissant, Guy Schofield, and Cade McCall. Emotion design for video games: a framework for affective interactivity. Games: Research and Practice, 1(3):19:1–19:??, September 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3624537.

Gomez:2023:EPI

[20] Argenis Ramirez Gomez and Michael Lankes. Eudaimonic player interaction: a reflection on input diegesis and meaning. Games: Research and Practice, 1(3):20:1–20:??, September 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3617670.

Pappas:2023:SSG

[21] Mike Pappas. Social safety in games: Moderating voice chat in the metaverse. Games: Research and Practice, 1(3):21:1–21:??, September 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3624356.

Davidson:2023:CSN

[22] Guy Davidson. The C++ standard needs to broaden its reach. Games: Research and Practice, 1(3):22:1–22:??, September 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3626207.

Deedman:2023:EOM

[23] Jon Deedman. Extremism online: Meaningful transparency for the gaming industry. Games: Research and Practice, 1(4):23:1–23:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3630026.

Frommel:2023:ICE

[24] Julian Frommel and Regan L. Mandryk. Individual control over exposure to combat toxicity in games. Games: Research and Practice, 1(4):24:1– 24:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3633768.

Xiao:2023:SAL

[25] Leon Y. Xiao. Shopping around for loot box presence warning labels: Unsatisfactory compliance on Epic, Nintendo, Sony, and Microsoft platforms. Games: Research and Practice, 1(4):25:1–25:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.

1145/3630631.

Grace:2023:GGE

[26] Lindsay D. Grace. Global games education now. Games: Research and Practice, 1(4):26:1–26:??, December 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3632364.

Tiwari:2023:ORG

[27] Siddhartha Paul Tiwari and Lindsay D. Grace. Observations and reflections on the gaming landscape in South East Asia. Games: Research and Practice, 1(4):27:1–27:??, December 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3632286.

Farough:2024:GIC

[28] Amanda Farough and Michael Futter. The game industry’s cataclysmic layoffs: Where do we go from here? Games: Research and Practice, 2(1): 1:1–1:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3651280.

Schlegel:2024:EAD

[29] Linda Schlegel. Extremist activities in digital gaming spaces: Challenges and opportunities for counter-extremism. Games: Research and Practice, 2 (1):2:1–2:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3649810.

White:2024:NGE

[30] Jessica White and Galen Lamphere-Englund. At the nexus of gaming and extremism. Games: Research and Practice, 2(1):3:1–3:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3643805.

Rusch:2024:CCT

[31] Doris C. Rusch. Co-creating transformative futures through the lens of deep game design. Games: Research and Practice, 2(1):4:1–4:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3643026.

Dunlap:2024:YBW

[32] Kelli N. Dunlap. You’ve (not) been warned: Developer best practices for content warnings and why compassion matters. Games: Research and Practice, 2(1):5:1–5:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic).

URL https://dl.acm.org/doi/10.1145/3638287.

Dunham:2024:PGA

[33] John Dunham, Jiangnan Xu, Konstantinos Papangelis, Nicolas Lalone, Michael Saker, and David Schwartz. Pokémon GO as an advertising platform: The case for locative advertising in location-based games. Games: Research and Practice, 2(1):6:1–6:??, March 2024. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3641509.

Hodent:2024:EGT

[34] Celia Hodent, Fran Blumberg, and Sebastian Deterding. Ethical games: Toward evidence-based guidance for safeguarding players and developers. Games: Research and Practice, 2(2):7:1–7:??, June 2024. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3685207.

Colvert:2024:PDE

[35] Angela Colvert, Kruakae Pothong, and Sonia Livingstone. Playful by design: Embedding children’s rights into the digital world. Games: Research and Practice, 2(2):8:1–8:??, June 2024. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3678469.

Stamou:2024:MIV

[36] Stefania Stamou, Konstantinos C. Apostolakis, Stavroula Ntoa, George Margetis, and Constantine Stephanidis. Museum-inspired video games as a symbolic transitional justice policy: Overview, concepts, and research directions. Games: Research and Practice, 2(2):9:1–9:??, June 2024. CO-DEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/

10.1145/3651279.

Reay:2024:TST

[37] Emma Reay. Treating symptoms or treating causes? Therapeutic video games for mental health. Games: Research and Practice, 2(2):10:1–10:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.

acm.org/doi/10.1145/3664282.

Kilmer:2024:AAC

[38] Elizabeth Kilmer, Zeynep Aslan, and Rachel Kowert. From avoidance to action: a call for open dialogue on hate, harassment, and extremism in the gaming industry. Games: Research and Practice, 2(2):11:1–11:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3656559.

Cho:2024:UDH

[39] Selina Cho, Jonathan Lusthaus, and Ivan Flechais. Unpacking the dynamics of harm in game cheating communities: a guiding framework for crossindustry intervention. Games: Research and Practice, 2(2):12:1–12:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3656558.

Chen:2024:TCC

[40] Sarah Chen. The toxic cost of cheap usernames: Re-applying Friedman and Resnick’s framework in competitive video game networks. Games: Research and Practice, 2(2):13:1–13:??, June 2024. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3669936.

Yang:2024:GHS

[41] Zachary Yang, Nicolas Grenon-Godbout, and Reihaneh Rabbany. Game on, hate off: a study of toxicity in online multiplayer environments. Games: Research and Practice, 2(2):14:1–14:??, June 2024. CODEN ???? ISSN 2832-

5516 (electronic). URL https://dl.acm.org/doi/10.1145/3675805.

Denoo:2024:MFM

[42] Maarten Denoo, Bruno Dupont, and Bieke Zaman. Many faces, many names? Ethics in Belgian game development education. Games: Research and Practice, 2(2):15:1–15:??, June 2024. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3675804.

Birk:2024:BDV

[43] Max V. Birk, Simone van der Hof, and Antonius J. van Rooij. Behavioral design in video games. Games: Research and Practice, 2(2):16:1–16:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.

acm.org/doi/10.1145/3672088.

Canca:2024:WGI

[44] Cansu Canca, Laura Haaber Ihle, and Annika Marie Schoene. Why the gaming industry needs responsible AI. Games: Research and Practice, 2 (2):17:1–17:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic).

URL https://dl.acm.org/doi/10.1145/3675803.

Petrovskaya:2024:APP

[45] Elena Petrovskaya. Ask the players! Player-centric principles as guidelines for ethical microtransactions. Games: Research and Practice, 2(2):18:1– 18:??, June 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https:

//dl.acm.org/doi/10.1145/3675802.

Blumberg:2024:AEG

[46] Fran C. Blumberg. Advancing ethics in the games industry: a round table. Games: Research and Practice, 2(2):19:1–19:??, June 2024. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3677008.